# NewsInterview: a Dataset and a Playground to Evaluate LLMs’ Grounding Gap via Informational Interviews

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive capabilities in generating coherent text but often struggle with grounding language and strategic dialogue. To address this gap, we focus on journalistic interviews, a domain rich in grounding communication and abundant in data. We curate a dataset of 40,000 two-person informational interviews from NPR and CNN, and reveal that LLMs are significantly less likely than human interviewers to use acknowledgements and to pivot to higher-level questions. Realizing that a fundamental deficit exists in multi-turn planning and strategic thinking, we develop a realistic simulated environment, incorporating source personas and persuasive elements, in order to facilitate the development of agents with longer-horizon rewards. Our experiments show that while source LLMs mimic human behavior in information sharing, interviewer LLMs struggle with recognizing when questions are answered and engaging persuasively, leading to suboptimal information extraction across model size and capability. These findings underscore the need for enhancing LLMs’ strategic dialogue capabilities.
@@
Co-first author contributions: Michael Lu did much of the implementation for discourse schemes and game design, as well as ideating the game idea. Alexander Spangher was the primary writer, he was the advisor for Michael and Sriya, and he performed the analysis in Section 4.2.
@@ 1 Introduction
Recent research has shown that LLMs struggle to engage in emotional Shaikh et al. (2024) or strategic Wongkamjan et al. (2024) dialogue. For example, Shaikh et al. (2024) examined LLM-generated responses to dialogues and found fewer occurrences of “grounding language” Clark (1996); cho-may-2020-grounding, like acknowledgements or affirmations, that humans typically use to foster comfort and trust. This can impede an LLM’s ability to serve in a variety of situations: e.g. education Kasneci et al. (2023), mental health Carlbring et al. (2023) or conflict resolution Argyle et al. (2023). However, prior efforts to ameliorate such gaps face limitations: existing large datasets (1k-10k transcripts) are generated via crowdsourcing and are inherently unnatural Rashkin et al. (2019); Wang et al. (2019); Liu et al. (2021). More natural transcripts, of educational Caines et al. (2020) or therapeutic environments Gratch et al. (2014), are difficult to collect due to privacy concerns Casey (2004) and are small-scale (100-1k transcripts).
@@
In this work, we directly address these limitations by focusing on an area where grounding communication is required but plentiful data exist: journalist interviews. Journalistic, or informational interviews, are typically conducted between an “interviewer” and a “source”, and the goal is to obtain information. Sources are often anxious or unclear Harcup (2015), and human interviewers are constantly evaluating: (1) Are my questions getting fully addressed? (2) Do I need to more effectively engage or persuade a source Sedorkin (2015)? This makes news interviews an ideal setting to observe grounding, strategic, and persuasive dialogues.
@@
To study how LLMs perform in journalistic contexts, we start by collecting interview transcripts from two major news sources: National Public Radio (NPR) and Cable News Network (CNN), filtering to over 40,000 dyadic informational interviews.We release our dataset and code at https://github.com/alex2awesome/news-interview-question-generation. Next, we show that LLMs in news interview settings suffer from the same lack of grounding as in other dialogue settings Shaikh et al. (2024). We find that significant discourse differences exist in the kinds of questions asked by LLMs and human interviewers: for example, LLMs are 50% less likely to make acknowledgements, and 30% less likely to pivot back to higher-level questions than humans.
@@
Next, we turn to a more fundamental question: why grounding? According to Cialdini (2009), grounding exists for a purpose: to influence an outcome (e.g. in therapeutic environments, the grounded patient is more open and makes more progress Bohart (1999); in educational environments, the comfortable student is more engaged and learns more Brown (1994); these are all usually borne out over many conversational turns). Observing the effects of grounding language in terms of objective outcomes might be a more effective way to reason and ultimately train empathetic agents and improve long-horizon strategic dialogue. Motivated by these observations, we develop a realistic game environment to serve as a playground: in this simulation, LLMs play the role of the interviewer and the source. The goal for the interviewer is to obtain the maximal amount of information from the source in a limited number of questions.
[…]
@@
  • We release a high-quality dataset of 40,000 two-person informational interviews from NPR and CNN. This dataset addresses the scarcity of large-scale dialogue data necessary for studying grounding communication.
@@ 2 Dataset Processing
In this section, we describe the collection and processing of our dataset.
(a) % of Discourse types throughout human interviews. Human journalists use different discourse roles across the interview, including gradually more Acknowledging statements, increasing from 5% at the start to over 20% by the end.
(b) % of Discourse types of LLM responses in interviews. LLMs display an increasing likelihood of asking opinion or broadening questions over the course of an interview and a lower likelihood of returning to outline-level questions.
Figure 1: Comparison of discourse types across interviews (turn 1, usually a greeting, is excluded). LLM is shown the first $k-1$ turns of a human interview and asked to generate the next question.
@@ 2.1 Data Collection
We aggregate, clean and condense multiple publicly available datasets of interview transcripts from National Public Radio (NPR) and Cable News Network (CNN) in order to build a high-quality interview dataset of 45k source-to-interview transcripts. These transcripts are published records of live interviews conducted between a journalist and sources invited on the program. They provide a rich resource for analyzing natural language interactions in informational interviews.
@@ 2.2 Data Filtering for Interview Analysis
We want to focus on one-on-one informational interviews between a journalist and a single source. We start with 487,310 transcripts collected by Majumder et al. (2020); Zhu et al. (2021). However, initial examination of the transcripts reveals many of them to be low-quality: many include multiple sources, are formatted as panel discussions, or are not informational in nature (e.g., they included game shows). To filter the transcripts and retain only those that fit our criteria, we prompt Llama-3.1-70bhttps://huggingface.co/meta-llama/Meta-Llama-3.1-70B-Instruct Touvron et al. (2023) using the vLLM framework Kwon et al. (2023) to classify each transcript based on the number of participants and the nature of the content. The prompts used for filtering are provided in App. C.5. After filtering, 45,848 interviews remain. Finally, the original transcripts do not distinguish which participant was the interviewer vs. the interviewee. So, we count each participant’s use of question marks: the participant with more is labeled the interviewer.Manual validation on 50 interviews showed this method correctly identified roles in $>98\%$ of cases.
@@
Table 1: Alignment of LLM-Generated Questions with Original Interview questions. We give an LLM, Llama-3.1-70b, the prior $k-1$ turns in an interview and prompt it to ask the next question. We measure the percentage of times this question aligns to a question asked by a human at the same point in the interview across six dimensions: Exact Match (nearly exactly the same as the original utterance), Information (relevant factual content), Motivation (same motivation as the original question), Style (alignment with tone and phrasing), Discourse (structural role within the interview), and Context (incorporation of contextual knowledge). The prompting strategies compared are Baseline-LLM, Chain-of-Thought (CoT), LLM with an Outline, and Outline-CoT; and, we conduct a human baseline trial with a former professional journalist.
Figure 2: Distribution of Discourse Roles in Questions, Across Different Prompting Strategies. We compare the proportions of discourse roles of questions (e.g. “Follow-up”, “Acknowledgement”, etc.) generated by (a) human journalists, (b) Baseline-LLM (Llama-3.1-70b) (c) LLM prompted with an Outline and (d) with Chain-of-Thought (CoT). Acknowledgement statements, which often build empathy, are significantly underrepresented in all LLM prompting approaches, compared to human-generated questions (see appendix for Outline-CoT).
@@ 3 Analysis
In this section, we analyze how humans conduct informational interviews and compare this behavior to that of pretrained LLMs. Our aim is to explore whether LLMs face similar grounding problems in interviews as observed in educational and therapeutic settings Clark (1996); shaikh-etal-2024-grounding.
@@ 3.1 Generating Counterfactual Utterances
One way to assess how an LLM would behave in an interview setting offline is to perform a counterfactual simulation shaikh-etal-2024-grounding. Specifically, given a human interview consisting of $n$ interviewer-source conversational turns $(q_{1},a_{1})...(q_{n},a_{n})$, we feed $t-1$ turns into the LLM along with a prompt instructing the LLM to generate the next question. This generates a counterfactual, $g_{t}$ to what the human would have said, $h_{t}$. We experiment with different variations: (1) Baseline: The LLM is simply asked to produce the next question. (2) Chain-of-Thought: The LLM is instructed to reason about the information already provided in the interview, consider what might be left to ask, and then generate the next question. (3) Outline: In addition to a chain-of-thought prompt, the LLM is also provided with an outline of the interview goals (described in Section 4.2) to incorporate into its chain-of-thought reasoning.We include full prompt examples for all three variations in Appendix C.5. All question-generation experiments are conducted using Llama-3.1-70b.
Figure 3: Walkthrough of the LLM Interviewer-Agent Process. To set up the interview, the interviewer agent is given a set of high-level objectives, similar to a journalist’s pre-interview notes, while the source is given a persona and a set of relevant facts. In each turn, the interviewer asks a question (Step 1). The source determines what information to reveal based on relevance and comfort level (Step 2a). Depending on the source’s comfort, a subset of relevant information is randomly selected for the response (Step 2b). The source then crafts a reply aligned with their persona (Step 2c). The reward given to the interview agent, at the end of $k$ turns, is the number of information items extracted from the source.
@@ 3.2 Evaluating LLM Counterfactuals
To analyze how similar LLM questions, $g_{t}$ are to human questions, $h_{t}$, we perform two analyses:
@@
We aim to assess how similar $g_{t}$ is to $h_{t}$ across different comparison categories Saha et al. (2024), specifically: Informational consistency (i.e. $g_{t}$ and $h_{t}$ seek similar informational objectives); Motivational, (i.e. similar outcomes); Style, (i.e. similar tone); Contextual consistency (i.e. similar appropriateness given the context); Discourse consistency (i.e. similar purposes in the overall conversation). Putting these together, we assess an $Exact$ match. We ask an LLM, GPT-4o, to perform this assessment and manually inspect its outputs and reasoning threads.
@@
[…]
@@ 3.3 Findings
As shown in Figure 2, grounding gaps exist in journalistic interviewing similar to those observed by Shaikh et al. (2024). While human journalistic interviewers tend to make Acknowledgement statements in about 9% of their utterances, all prompting variations that we experimented with made close to zero acknowledging statements. This lack of acknowledgement is paired with not mirroring the source’s speaking style; human journalists, as shown in Appendix C.6, bring character and voice.
Insight #2: LLMs do not engage in strategic multi-turn questioning
. Even in settings where LLMs are exposed to interview outlines, they are still undirected in their questions. As shown in Figure 2, LLMs are significantly more likely to ask follow-up questions than humans across all prompting variations. Introducing chain-of-thought and outline variations increases the rate at which the LLM asks outline-level questions. However, the rate remains significantly below human levels. Additionally, they are also more likely to ask either Opinion questions or Broadening questions. In fact, in Figure 1(b), we observe that LLMs tend to ask increasing amounts of Opinion Questions and Broadening Questions over time, which humans do not. As can be seen in Table 8, these questions can be vague and open-ended and reflect a lack of planning. Together, these findings suggest an inability to direct an interview in a desired direction and engage in multi-turn planning.
Insight #3: LLMs are capable of understanding context, but fail in all other categories of similarity to humans.
Comparing the content and style of LLM interviews to human interviews in Table 1, we note that, overall, LLMs are broadly dissimilar to humans in style, motivation and information-seeking. One area where the LLMs succeed, relatively, is understanding the context of the interview beforehand. This is not a new observation – much recent work, e.g. in dialogue-tracking, has found LLMs to perform well Ou et al. (2024). The fact that LLMs can preserve context over multiple turns and do not drift away from the topic indicates that models might one day be able to engage in multi-turn goal-oriented dialogue, given the right reward signals and learning environment.
@@
4 NewsInterview: An Interview Game
As shown, LLM counterfactual questions exhibit several shortcomings: they are less likely to acknowledge the interviewee and focus excessively on follow-up questions. However, questions remain: do they lack strategic multi-turn planning? In human dialogue, grounding exists ultimately strategic purposes Cialdini (2009). Our goal for the remainder of the paper is to set up a realistic simulated game-environment with a delayed reward signal to test this.
4.1 Game Design Overview
We first introduce our game on a high level, explain our game design, as illustrated in Figure 3, and then describe our implementation. We draw on two journalism textbooks: Interviewing: A Guide for Journalists and Writers, which explains how to conduct effective interviews and speak to reluctant, defensive, or poor-explaining sources Sedorkin (2015); and Journalism: Principles and Practice, which describes how to build source trust Harcup (2015).
Our gameplay proceeds in a loop, shown in Algorithm 1. The “player” in our game plays the role of an interviewer and is able to ask questions to a source, based on the conversational history and the interview objectives (the Interviewer() step). The source is given a set of informational items, and assesses whether any of these items are relevant to the question (the getRelevantInfo() step); they then decide how persuaded or comfortable they are based on the conversational history (the getPersuasionLevel() step); based on this, we determine the subset of relevant items they return (the getItemsToReturn()). They respond with these items. The reward, obtained at the end of the game, is that amount of information items the source disclosed.
4.2 Gameplay Design
We first start by describing our data processing, and then we will describe the functions introduced in Algorithm 1 in more detail.
@@ 4.2 Gameplay Design
3$\triangleright$ Step 1: Interviewer Question Generation
4:     $q=$ Interviewer$(c,o)$
5$\triangleright$ Step 2: Source’s Response Generation
6:     $r=$getRelevantInfoItems$(I,u,q)$
7:     $p=$getPersuasionLevel$(c)$
8:     $f=$getItemsToReturn$(r,p)$
9:     $a=$Source$(q,c,f,p,\phi)$
10$\triangleright$ Update Variables
11:     $u\leftarrow u\cup f$, $c\leftarrow c\oplus\left[q,a\right]$, $R\leftarrow R+|f|$

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive capabilities in generating coherent text but often struggle with strategic dialogue. To address this gap, we focus on journalistic interviews. We curate a dataset of 40,000 two-person informational interviews from major news organizations in scenarios where human interviewers employ strategies to coax information from sources. We then try to mimic these activities with LLMs and find striking differences; models are less likely to use acknowledgments and more likely to rabbit-hole and not pivot to other topics. Realizing that a fundamental deficit exists in LLM multi-turn planning and strategic thinking, we develop a realistic simulated environment, incorporating source personas and persuasive elements, in order to facilitate the development of agents with long-horizon rewards. Our experiments show that mimicry failures are not two-sided; when posing as a source, models adequately reflect human behavior in information sharing, making our simulation a realistic benchmark. Interviewer-LLMs, however, struggle with engaging persuasively, leading to suboptimal information extraction across model size and capability. This simulated game lays the groundwork for future work in enhancing LLMs’ strategic dialogue capabilities. [^1] We release dataset and code at https://github.com/alex2awesome/news-interview-question-generation.
@@
(a) Proportion of Discourse types throughout human interviews. Human journalists use different discourse roles across the interview, including gradually more Acknowledging statements, increasing from 5% at the start to over 20% by the end.
(b) Proportion of Discourse types of LLM responses in interviews. LLMs display an increasing likelihood of asking opinion or broadening questions over the course of an interview and a lower likelihood of returning to outline-level questions.
Figure 1: Comparison of discourse types across interviews (the first turn, usually a greeting, is excluded). The LLM is shown the first t1t-1 turns of a human interview and asked to generate the next question.
@@ 1 Introduction
Recent research has shown that LLMs struggle to engage in emotional Shaikh et al. (2024a) or strategic Wongkamjan et al. (2024) dialogue. For example, Shaikh et al. (2024a) examined LLM-generated responses to dialogues and found fewer occurrences of “grounding language” Clark (1996); Cho and May (2020), like acknowledgements or affirmations, that humans typically use to foster comfort and trust. This can impede an LLM’s ability to serve in a variety of situations: e.g., education Kasneci et al. (2023), mental health Carlbring et al. (2023) or conflict resolution Argyle et al. (2023). However, prior efforts to ameliorate such gaps face limitations: existing large datasets (1k–10k transcripts) are generated via crowdsourcing and are inherently unnatural Rashkin et al. (2019); Wang et al. (2019); Liu et al. (2021). More natural datasets, of educational Caines et al. (2020) or therapeutic environments Gratch et al. (2014), are difficult to collect due to privacy concerns Casey (2004) and are small-scale (100–1k transcripts).
@@
In this work, we directly address these limitations by focusing on an area where grounding communication is required but plentiful data exist: journalist interviews. Journalistic, or informational interviews, are typically conducted between an “interviewer” and a “source,” and the goal is to obtain information. Sources are often anxious or unclear Harcup (2015), and human interviewers are constantly evaluating: (1) Are my questions getting fully addressed? (2) Do I need to more effectively engage or persuade a source Sedorkin (2015)? This makes news interviews an ideal setting to observe grounding dialogues.
@@
To study how LLMs perform in journalistic contexts, we start by collecting interview transcripts from two major US news sources: National Public Radio (NPR) and Cable News Network (CNN), filtering to over 40,000 dyadic informational interviews. [^2] As opposed to games, questionaires and other formats these news outlets release Next, we show that LLMs in news interview settings suffer from the same lack of grounding as in other dialogue settings Shaikh et al. (2024a). We find that significant discourse differences exist in the kinds of questions asked by LLMs: for example, LLMs are 50% less likely to make acknowledgements, and 30% less likely to pivot to higher-level questions.
@@
Next, we turn to a more fundamental question: why consider grounding? According to Cialdini (2009), grounding exists for a purpose: to influence an outcome (e.g., in therapeutic environments, the grounded patient is more open and makes more progress Bohart and Tallman (1999); in educational environments, the comfortable student is more engaged and learns more Brown and Campione (1994); these are all usually borne out over many conversational turns). Observing the effects of grounding language in terms of objective outcomes might be a more effective way to reason and ultimately train empathetic agents and improve long-horizon strategic dialogue. Motivated by these observations, we develop a realistic game environment to serve as a playground: in this simulation, LLMs play the role of the interviewer and the source. The goal for the interviewer is to obtain the maximal amount of information from the source in a limited number of questions.
In order to induce the need for grounding communication, we design different personas for sources (e.g., anxious, clueless, dominating), each with different communication patterns. We also add a responsiveness to strategic dialogue: sources will only return information if they are persuaded in a manner befitting their personas [^3] We understand that “being persuaded,” “being made comfortable,” and “being acknowledged” are all separate forms of grounding, some more active than others. However, we use “persuasion” as a short-hand encompassing all categories. Harcup (2015); Sedorkin (2015). We find that our environment is realistic: source-LLMs correlate significantly with humans in their ability to identify persuasion ($r=.43$, $p<.0001$). However, interviewer-LLMs struggle to both recognize when questions are answered and actively persuade the source, resulting in suboptimal information extraction.
@@
  • We release a high-quality dataset of 40,000 dyadic informational interviews from NPR and CNN. This dataset addresses the scarcity of large-scale dialogue data necessary for studying grounding communication.
@@ 2.1 Data Collection
We aggregate, clean and condense multiple publicly available datasets of interview transcripts from NPR and CNN in order to build a high-quality interview dataset of 45k source-to-interview transcripts. These transcripts are published records of live interviews conducted between a journalist and sources invited on the program. They provide a rich resource for analyzing natural language interactions.
@@ 2.2 Data Filtering for Interview Analysis
We want to focus on one-on-one informational interviews between a journalist and a single source. We start with 487,310 transcripts collected by Majumder et al. (2020) and Zhu et al. (2021). However, initial examination of the transcripts reveals many of them to be low-quality: they include multiple sources, are formatted as panel discussions, or are not informational in nature (e.g., they include game shows). To filter the transcripts and retain only those that fit our criteria, we prompt Llama-3.1-70b [^4] https://huggingface.co/meta-llama/Meta-Llama-3.1-70B-Instruct Touvron et al. (2023) using the vLLM framework Kwon et al. (2023) to classify each transcript based on the number of participants and the nature of the content. The prompts used for filtering are provided in App. B.5. After filtering, 45,848 interviews remain. Finally, the original transcripts do not distinguish which participant was the interviewer vs. the interviewee. So, we count each participant’s use of question marks: the participant with more is labeled the interviewer. [^5] Manual validation on 50 interviews showed this method correctly identified roles in $98\%$ of cases.
Conversations have, on average, 7.5 turns between the interviewer and source. The source speaks for longer, with an average of 551 words per conversation compared with the interviewer’s 270 words (or 27 words per source utterance, 16 per interviewer). Interviewers tend to ask “what” and “how” questions the most, and conversations occur at Flesch-Kincaid Grade of 6.9 Kincaid et al. (1975). Interviews cover a range of topics, from literature, politics, academics, and international affairs (see Appendix A for more details).
@@
Table 1: Alignment of LLM-Generated Questions with Original Interview questions. We give an LLM, Llama-3.1-70b, the prior t1t-1 turns in an interview and prompt it to ask the next question. We measure the percentage of times this question aligns to a question asked by a human at the same point in the interview across six dimensions: Exact Match (nearly exactly the same as the original utterance), Information (relevant factual content), Motivation (same motivation as the original question), Style (alignment with tone and phrasing), Discourse (structural role within the interview), and Context (incorporation of contextual knowledge). The prompting strategies compared are Baseline-LLM, Chain-of-Thought (CoT), LLM with an Outline, and Outline-CoT; and, we conduct a human baseline trial with a former professional journalist.
Figure 2: Distribution of Discourse Roles in Questions, Across Different Prompting Strategies. We compare the proportions of discourse roles of questions (e.g., “Follow-up,” “Acknowledgement,”) generated by (a) human journalists, (b) Baseline-LLM (Llama-3.1-70b) (c) LLM prompted with an Outline and (d) with Chain-of-Thought (CoT). Acknowledgement statements, which often build empathy, are significantly underrepresented in all LLM prompting approaches, compared to human-generated questions (see appendix for Outline-CoT).
@@ 3 Analysis
In this section, we analyze how humans conduct informational interviews and compare this behavior to that of pretrained LLMs, to explore whether LLMs face similar grounding problems as observed in other settings Clark (1996); Shaikh et al. (2024b).
@@ 3.1 Generating Counterfactual Utterances
One way to assess how an LLM would behave in an interview setting offline is to perform a counterfactual simulation Shaikh et al. (2024b). Specifically, given a human interview consisting of at least $t$ interviewer-source conversational turns $(q_{1},a_{1})...(q_{t},a_{t})$..., we feed t1t-1 turns into the LLM along with a prompt instructing the LLM to generate the next question. This generates a counterfactual, $g_{t}$ to what the human would have said, $q_{t}$. We experiment with different variations: (1) Baseline: The LLM is simply asked to produce the next question. (2) Chain-of-Thought (CoT): The LLM is instructed to reason about the information already provided in the interview, consider what might be left to ask, and then generate the next question. (3) Outline: the LLM is provided with an outline of the interview goals (described in Section 4.2) to incorporate into CoT reasoning. [^6] We include full prompt examples for all three variations in Appendix B.5. All question-generation experiments are conducted using Llama-3.1-70b.
@@ 3.2 Evaluating LLM Counterfactuals
To analyze how similar LLM questions, $g_{t}$ are to human questions, $q_{t}$, we perform two analyses:
@@
We aim to assess how similar $g_{t}$ is to $q_{t}$ across different comparison categories Saha et al. (2024), specifically: Informational consistency (i.e., $g_{t}$ and $q_{t}$ seek similar informational objectives); Motivational, (i.e., similar outcomes); Style, (i.e., similar tone); Contextual consistency (i.e., similar appropriateness given the context); Discourse consistency (i.e., similar purposes in the overall conversation). Putting these together, we assess an $Exact$ match. We ask an LLM, GPT-4o, to perform this assessment and manually inspect its outputs and reasoning threads.
@@
We aim to assess whether $g_{t}$ plays a similar function as $q_{t}$ does. We develop a schema to describe the role of each question. [^7] To generate our discourse schema, we asked two journalists to analyze fifty interview transcripts. One had eight years of experience in newsrooms, the other was an undergraduate student studying journalism. We held three conferencing sessions to develop the schema. Then, we blindly annotated ten interviews, achieving a $\kappa=.6$. Given our schema, we then asked an LLM to classify discourse roles in sentences. The prompt contains the interview context, $(q_{1},a_{1})...(q_{t-1},a_{t-1})$, and current question $q_{t}$. To validate the LLM’s labeling accuracy, we had the professional journalist label 10 additional interviews as ground-truth and scored the LLM’s assignments. The LLM scored a .8.8 f1 score. This schema includes the following elements: Follow-up Question (e.g., “Can you tell us more?”), Outline-Level Question (e.g., “Moving on, can we discuss the next event?”), Acknowledgement Statement (e.g., “I see, that sounds scary.”), Opinion/Speculation (e.g., “What do you think will happen?”), Broadening Question (e.g., “How does this fit into the broader trend?”), Verification Question (e.g., “So to confirm…”) and Challenge Question (e.g., “These dates don’t line up.”). See Table 13 in the Appendix for definitions of each role.
@@ 3.3 Findings
As shown in Figure 2, grounding gaps exist in journalistic interviewing similar to those observed by Shaikh et al. (2024a). While human journalistic interviewers tend to make Acknowledgement statements in about 9% of their utterances, all prompting variations that we experimented with made close to zero of these statements. This lack of acknowledgement is paired with not mirroring the source’s speaking style; human journalists, as shown in Appendix B.6, bring character and voice.
Insight #2: LLMs do not engage in strategic multi-turn questioning.
Even in settings where LLMs are exposed to interview outlines, they are still undirected in their questions. As shown in Figure 2, LLMs are significantly more likely to ask follow-up questions than humans across all prompting variations. Introducing chain-of-thought and outline variations increases the rate at which the LLM asks outline-level questions. However, the rate remains significantly below human levels. Additionally, they are also more likely to ask either Opinion questions or Broadening questions. In fact, in Figure 1(b), we observe that LLMs tend to ask increasing amounts of Opinion Questions and Broadening Questions over time, which humans do not. As shown in Table 8, these questions can be vague and open-ended. Together, these findings suggest an inability to direct an interview in a desired direction and engage in multi-turn planning.
Insight #3: LLMs are capable of understanding context, but fail in other categories of similarity to humans.
Comparing the content and style of LLM interviews to human interviews in Table 1, we note that, overall, LLMs are broadly dissimilar to humans in style, motivation and information-seeking. One area where the LLMs succeed, relatively, is understanding the context of the interview beforehand. This is not a new observation – much recent work, e.g., in dialogue-tracking, has found LLMs to perform well Ou et al. (2024). The fact that LLMs can preserve context over multiple turns and do not drift away from the topic indicates that models might one day be able to engage in multi-turn goal-oriented dialogue, given the right reward signals and learning environment.
@@
[…]
4:  $g_{i}=$ Interviewer(C,o)(C,o)
[…]
7:  p=p=getPersuasionLevel(C)(C)
8:  f=f=getItemsToReturn(r,p)(r,p)
[…]
11:  $U\leftarrow U\cup f$, $C\leftarrow C\oplus\left[g_{i},a_{i}\right]$, $R\leftarrow R+|f|$
@@
Figure 3: Walkthrough of the LLM Interviewer-Agent Process. In the NewsInterview game, an interviewer-LLM converses with a source-LLM: the interviewer-LLM is rewarded based on how many information items (shown at the right) are extracted from the source. In more detail: the interviewer agent is given a set of high-level objectives, similar to a journalist’s pre-interview notes, while the source is given a persona and a set of relevant facts. The interview proceeds for $k$ turns. Interviewer Query: the interviewer is prompted to ask a question based on their goals and information obtained (Step 1). Source Response: The source responds with a multi-step process. First, they are prompted to determine how many information items in their factbook are relevant to the question (Step 2a). Then, they self-assess their comfort level. Depending on this, the simulation randomly selects a subset of relevant information for the response (Step 2b). We track the decision of which items to return on the back-end, in order to calculate the final reward. The source is then prompted to craft a reply aligned with their persona (Step 2c). After $k$ turns: a reward given to the interviewer based on the number of information items extracted from the source.
4 NewsInterview: An Interview Game
As shown, LLM counterfactual questions exhibit several shortcomings: they are less likely to acknowledge the interviewee and focus excessively on follow-up questions. But do both of these shortcomings point to a lack of strategic multi-turn planning? In human dialogue, grounding exists for long-term strategic purposes Cialdini (2009), yet there currently exists no way to way to obtain these kinds of long-term rewards during LLM training. Motivated by this insight, our goal for the remainder of the paper is to create and validate a realistic game-environment with a delayed reward signal. We leave to future work utilization of this framework for improving strategic dialogue.
4.1 Game Design Overview
We first introduce our game on a high level, illustrated in Figure 3, and then describe our implementation. Our gameplay proceeds in a loop, shown in Algorithm 1. The “player” in our game plays the role of an interviewer and is able to ask questions to a source, based on the conversational history and the interview objectives (the Interviewer() step). The source is given a set of informational items and assesses whether any of these items are relevant to the question (the getRelevantInfoItems() step); the source then decides how persuaded or comfortable they are based on the conversational history (the getPersuasionLevel() step). Based on this, we determine the subset of relevant items the source returns (the getItemsToReturn()), and track these on the back-end as an accumulating reward. The reward, obtained at the end of the game, is the unique number of information items disclosed.
4.2 Gameplay Design
To design our game, we draw heavily on two journalism textbooks: Interviewing: A Guide for Journalists and Writers, which explains how to conduct effective interviews and speak to reluctant, defensive, or poor-explaining sources Sedorkin (2015); and Journalism: Principles and Practice, which describes how to build trust Harcup (2015). We first start by describing our data processing, and then we will describe Algorithm 1 in more detail. For all gameplay prompts, see Appendix C.